Circular No. 8797

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 B1 (CHRISTENSEN) E. J. Christensen reports his discovery of a comet on CCD images ob-

tained with the Catalina Sky Survey’s 0.68-m Schmidt telescope (discovery observation tabulated below); four stacked 30-s discovery images, taken in good seeing, show a slightly condensed 10′′ coma, while four stacked 90-s follow-up images reveal a 15′′ tail oriented toward p.a. 220^{o}–250^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several other CCD observers have noted the object’s cometary appearance, including G. Sostero and L. Donato at Remanzacco, Italy (0.45-m f/4.4 reflector, Jan. 20.04–20.09 UT; coma diameter 10′′ with a 16′′ tail toward p.a. 250^{o}); D. Mayes at Table Mountain (0.6-m reflector, Jan. 20.4); and L. Buzzi at Varese, Italy (0.60-m f/4.6 reflector, Jan. 20.0; 60 stacked 15-s exposures show a compact 10′′ coma with a possible elongation in p.a. ∼ 230^{o}).

2007 UT  α_{2000}  δ_{2000}  Mag.
Jan. 17.27421  8^{h}57^{m}29.^{s}84  +32^{o}29′41′′.0  18.6

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-B25.

T = 2007 Feb. 8.351 TT  ω = 54.^{o}294
e = 0.57317  Ω = 76.702 }2000.0
q = 2.42086 AU  i = 12.114
a = 5.67165 AU n^{o} = 0.072969 P = 13.5 years

COMET C/2006 P1 (McNAUGHT) Unfiltered-CCD total-magnitude estimates in daylight: Jan. 13.14 and

14.13 UT, −5.0 (A. Nakamura, Kuma, Ehime, Japan, 7.2-cm reflector; integrated coma diameter 1′.4); 15.48, −5.2 (K. Hornoch, J. Srba, and E. Brezina, Vsetin Observatory, Czech Republic, 6.3-cm reflector + neutraldensity filter; aperture size 8′). Further visual total-magnitude estimates: Jan. 16.94, −2.8 (S. J. O’Meara, Mauna Loa, HI, U.S.A., 10×50 binoculars, daylight; 0^{o}.5 tail); 18.50, −2.1 (A. Pearce, City Beach, W. Australia, naked eye; 4^{o}.5 tail); 20.51, −0.7 (Pearce; 23^{o} tail; tail synchrones easily visible).

COMET 183P/KORLEVIĆ-JURIĆ Comet P/2006 Y1 = 1999 DN_{3} (cf. IAUC 8786) has been numbered

183P (cf. MPC 58485).

2007 January 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green